A doctor measured serum HDL levels in her patients, and found that they were normally distributed with a mean of 64.9 and a standard deviation of 2.7. Find the serum HDL levels that correspond to the following z-scores. Round your answers to the nearest tenth, if necessary.

Z=-1.4

Answer:____

Z=2.36

Answer:____
Answer #1

Solution :

(a)

Z = -1.4

x = z * \sigma + \mu = -1.4 * 2.7 + 64.9 = 61.1

Answer = 61.1

(b)

Z = 2.36

x = z * \sigma + \mu = 2.36 * 2.7 + 64.9 = 71.3

Answer = 71.3